Science Report – November 21st

Science Report 21 November 2019
SOL: 11
Crew 215 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Dome Operations: The experimenting with gypsum and sand to create a form of plaster of paris cement was completed. After drying, the tiles will be ready for strength testing. Approximately 10ml of water was collected as a product of the heating and will be tested as suitable for medical solutions. In the afternoon, fine grained, pure gypsum was heated and prepared as a plaster of paris and a small splint was successfully molded and used to immobilize an index finger.
Green Hab: Green Hab operations have continued with additional seeds and seedlings being planted. See Green Hab report.
RAM Operations: N/A
EVA Suit Maintenance: Nominal. See Operations Report.
Psycho-social Stressors Monitoring: Data collection continues through the hexoskin(TM) monitoring vests. A final checklist-supported response simulated emergency was carried out in the afternoon to further test the stress on untrained personnel involved in medical emergencies in isolated situations.
Additional activities: N/A

Copyright © The Mars Society. All rights reserved. | Main Site